# Stopping rules for Monte Carlo methods of martingale difference type

### v1

@@ 5.2 Adaptive control variates
In this section, we examine the problem setting described in Examples 2.2 (c) and 3.1 (c). To recall it, suppose, for $\Psi:(0,1)^{d}\to\mathbb{R}$, we wish to evaluate the integral $\mu=\int_{(0,1)^{d}}\Psi({\bf u})d{\bf u}$, where the random element here is modeled via the uniform law on the unit hypercube $(0,1)^{d}$, which can be considered general enough thanks to the inverse transform random sampling method. Here, without affecting the value of the integral, one can introduce a parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ in a linear manner, as follows
@@
with ${\bm{\theta}}(0)=0_{d}$. By representing ${\bm{\theta}}(t)$ as a function of the batch index, we stress that it is updated only once for each batch. It clearly remains true that ${\bm{\theta}}(t)\to{\bm{\theta}}^{*}$, $\mathbb{P}_{0}$-$a.s.$ as $t\to+\infty$, just as in (5.6), because this is still nothing but standard iid Monte Carlo methods for the rightmost expression of (5.4). Here, the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ can be set as $\mathcal{F}_{k}=\sigma(\mathcal{N})\vee\sigma(U_{1},\cdots,U_{k})$ for $k\in\mathbb{N}$, as well as the branched one $(\mathcal{F}_{k-1}^{\star})_{k\in M(\tau)}$ by $\mathcal{F}_{k-1}^{\star}=\mathcal{F}_{m(\tau_{\epsilon,\delta})}\vee\sigma(U_{m(\tau_{\epsilon,\delta}-1)+1}^{\star},\cdots,U_{k-1}^{\star})$ for all $k\in M(\tau_{\epsilon,\delta})$ on the iid resampled batch $\{U_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, where $\tau_{\epsilon,\delta}$ is a suitable stopping batch as in (5.1). Then, for every $t\in\mathbb{N}$ and $k\in M(t)$, the base formulation (Assumption 2.1) holds by setting $X_{k}=\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle$, as
@@
It is worth noting that the proposed stopping rules for adaptive control variates have an interesting connection to averaged stochastic gradient descent (ASGD) [2]. While conventional SGD implementations often rely on validation performance for early stopping [28], monitoring averaged iterates, which helps reduce the effect of the noise term, has proven beneficial in practice. This suggests potential value in developing statistically driven stopping criteria that leverage averaging, as in our framework, particularly for validation-free scenarios [5]. ∎
@@ 6 Concluding remarks
We conclude this work by highlighting two promising directions for future research. The first involves improving the quality of the proposed stopping rules. Unlike stopping rules based on the classical central limit theorem (such as [1, 17]), a key challenge in enhancing the quality of stopping rules here arises from the lack of accurate finite-time approximations for the limiting law in the context of the martingale central limit theorem. Specifically, obtaining Berry-Esseen-type bounds based on higher-order moments, and moreover presenting them in an implementable form, remains out of reach. Although there is extensive literature on martingale central limit theorems and their rates of convergence - typically expressed as uniform bounds (such as [4, 7, 24]) and, in some cases, as non-uniform bounds (such as [14, 15, 19]) - these results generally involve unspecifiable constants, limiting their practical applicability to the development of stopping rules, with an exception [6], which, to the best of our knowledge, provides the only implementable result of the uniform bound of Berry-Esseen type.
@@ 7 Proofs
Proof of Theorems 2.4 and 3.2.
Theorem 2.4 is elementary by the tower rule and the standing assumption $\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the unbiasedness, as well as due to, for instance, [8, VII.9] for the strong law of large numbers. For Theorem 3.2, the first result can be derived as
@@
where we have applied the $\mathcal{F}_{m(\tau)}$-measurability of the stopping batch $\tau$ for the first equality and the tower rule with $\mathbb{E}^{\star}_{k-1}[X_{k}^{\star}]=\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the third and fourth equalities, due to $\mathcal{F}_{m(\tau-1)}\subseteq\mathcal{F}^{\star}_{k-1}$ for all $k\in M(\tau)=\{m(\tau-1)+1,\cdots,m(\tau)\}$. The second result is then straightforward due to $\mathcal{F}_{0}\subseteq\mathcal{F}_{m(\tau)}$. ∎
@@
A standard martingale central limit theorem (for instance, [16, Theorem 2.5 (a)] with $\xi_{t,k}=(X_{k}-\mu)/\sqrt{|M(t)|v^{2}_{0}(t)}$) reads in the present framework as follows. If $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ and $(|M(t)|v^{2}_{0}(t))^{-1}\sum_{k\in M(t)}\mathbb{E}_{k-1}[|X_{k}-\mu|^{2}\mathbbm{1}(|X_{k}-\mu|>\epsilon\sqrt{|M(t)|}v_{0}(t))]\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}0$ for every $\epsilon>0$, then it holds that $\mathbb{P}_{0}(\sqrt{|M(t)|}(\mu(t)-\mu)/v_{0}(t)\leq x)\to\Phi(x)$ for all $x\in\mathbb{R}$, as $t\to+\infty$. This Lindeberg condition can be simplified as the one in Theorem 4.2, thanks to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ and $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ (Assumption 4.1). The desired weak convergence (4.6) holds true by the Slutsky theorem due to the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$. ∎
@@
$\epsilon^{1/q}\tau_{\epsilon,\delta}^{0}$ $=\epsilon^{1/q}\mathop{\rm argmin}_{t\in\mathbb{N}}\left\{\mathbb{P}_{0}\left(\sqrt{|M(t)|}\frac{|\mu(t)-\mu|}{v_{0}(t)}>\sqrt{|M(t)|}\frac{\epsilon}{v_{0}(t)}\right)\leq\delta\right\}$
$\sim\mathop{\rm argmin}_{t\in(0,+\infty)}\left\{2\left(1-\Phi\left(\epsilon\sqrt{|M(\lfloor\epsilon^{-1/q}t\rfloor)|}/v(\lfloor\epsilon^{-1/q}t\rfloor)\right)\right)\leq\delta\right\},$
@@
For Proposition 4.5, the expression (4.7) can be rewritten as $\tau^{a}_{\epsilon,\delta}=\min_{t\in\mathbb{N}}\{\epsilon\sqrt{|M(t)|}/(v(t)+a(t))\geq\Phi^{-1}(1-\delta/2)\}$, which yields $\tau^{a}_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ as $\epsilon\to 0+$ or $\delta\to 0+$, due to $|M(t)|\to+\infty$ (Assumption 2.3), $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1). With the aid of this expression, we get $\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}/(v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta}))\geq\Phi^{-1}(1-\delta/2)$ at the time $m(\tau^{a}_{\epsilon,\delta})$, and $\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}/(v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1))<\Phi^{-1}(1-\delta/2)$ at the time $m(\tau^{a}_{\epsilon,\delta}-1)$, which together give, in a similar manner to (7.2),
$\Phi^{-1}(1-\delta/2)\leq\frac{\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}$
$\,=\frac{\epsilon\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}\frac{\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}\frac{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}$
$\,<\Phi^{-1}(1-\delta/2)\frac{\sqrt{|M(\tau^{a}_{\epsilon,\delta})|}}{\sqrt{|M(\tau^{a}_{\epsilon,\delta}-1)|}}\frac{v(\tau^{a}_{\epsilon,\delta}-1)+a(\tau^{a}_{\epsilon,\delta}-1)}{v(\tau^{a}_{\epsilon,\delta})+a(\tau^{a}_{\epsilon,\delta})}\stackrel{{\scriptstyle a.s.}}{{\to}}\Phi^{-1}(1-\delta/2),$ (7.4)
as $\epsilon\to 0+$, due to the first result $\tau^{a}_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$, as well as $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4), $v^{2}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1), and $\lim_{t\to+\infty}a(t)=0$. Hence, Part (ii) holds by the continuous mapping theorem since $\Phi$ is continuous on $\mathbb{R}$. With the additional condition, the first result of Part (iii) can be obtained by reformulating the equation (7.4). The second result of Part (iii) holds by the asymptotic equivalence (7.3) with the resampled empirical mean $\mu_{\star}$ and the stopping rule $\tau_{\epsilon,\delta}^{a}$, which yields the desired weak convergence due to the martingale central limit theorem (4.6) with the aid of $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\tau_{\epsilon,\delta}^{a}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ (Proposition 4.5 (i)), where the two asymptotic equivalences above hold almost surely, respectively, due to the first result and the additional condition of Part (iii). ∎
@@ A.1 ARCH(1)
Recall that $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$ for all $k\in\mathbb{N}$ with $X_{0}=0$ and $\{V_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with the $t$-distribution with $n$ degrees of freedom for some $n\in\{5,6,\cdots\}$, scaled in such a way to satisfy $\mathbb{E}_{0}[V_{1}^{2}]=1$, that is, the common probability density function is given by
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((n+1)/2)}{\sqrt{(n-2)\pi}\Gamma(n/2)}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2},\quad x\in\mathbb{R}.$ (A.1)
In particular, we have $\mathbb{E}_{0}[|V_{1}|^{4}]=3(n-2)/(n-4)$. It is trivial to derive that $\mathbb{E}_{k-1}[X_{k}]=(\beta+\alpha X_{k-1}^{2})^{1/2}\mathbb{E}_{k-1}[V_{k}]=0$ and $\mathbb{E}_{k-1}[X_{k}^{2}]=\beta+\alpha X_{k-1}^{2}$, due to $X_{k-1}\in\mathcal{F}_{k-1}$, $V_{k}\perp\mathcal{F}_{k-1}$, and $\mathbb{E}_{k-1}[V_{k}^{2}]={\rm Var}_{k-1}(V_{k})={\rm Var}_{0}(V_{k})=1$ for all $k\in\mathbb{N}$. Hence, we get $\mathbb{E}_{0}[X_{k}]=0$ (Assumption 2.1 with $\mu=0$) and ${\rm Var}_{0}(X_{k})=\mathbb{E}_{0}[X_{k}^{2}]=\beta+\alpha\mathbb{E}_{0}[X_{k-1}^{2}]=\cdots=\beta(1-\alpha^{k})/(1-\alpha)>0$, due to $\alpha\in(0,1)$, $\beta\in(0,+\infty)$ and $X_{0}=0$. Hence, the discrete-time stochastic process $\{X_{k}:\,k\in\mathbb{N}_{0}\}$ is asymptotically weakly stationary with mean $\mathbb{E}_{0}[X_{k}]=0$ and variance $\lim_{k\to+\infty}{\rm Var}_{0}(X_{k})=\beta/(1-\alpha)$, due to $\alpha\in(0,1)$. Hence, we get the convergence of the theoretical batch variance as
@@
One can show that all the relevant conditions hold true if the constant $\alpha$ and the degree of freedom $n$ satisfy $3\alpha^{2}(n-2)/(n-4)<1$. To verify the convergence of the conditional batch variance, observe first that for every $k\in\mathbb{N}$ and $m\in\{k+1,\cdots\}$,
@@
$=\frac{3(n-2)}{n-4}\left(\beta^{2}+2\beta\alpha\mathbb{E}_{0}\left[X_{k-1}^{2}\right]+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right)$
$\sim\frac{3(n-2)}{n-4}\left(\beta^{2}\frac{1+\alpha}{1-\alpha}+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right).$
If $3\alpha^{2}(n-2)/(n-4)<1$, then we get
$\mathbb{E}_{0}\left[|X_{k}|^{4}\right]\to\frac{3\beta^{2}(1+\alpha)(n-2)}{(1-\alpha)(n-4-3\alpha^{2}(n-2))},$
${\rm Var}_{0}\left(X_{k}^{2}\right)\to\frac{3\beta^{2}(1+\alpha)(n-2)}{(1-\alpha)(n-4-3\alpha^{2}(n-2))}-\frac{\beta^{2}}{(1-\alpha)^{2}},$
@@
$\mathbb{P}_{0}$-$a.s.$, as $t\to+\infty$. With $v^{2}_{0}(t)\to\beta/(1-\alpha)$, the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ (Theorem 4.4) holds true. In addition, the convergence of the sequence $\{\mathbb{E}_{0}[|X_{k}|^{4}]\}_{k\in\mathbb{N}}$ above implies the condition $\limsup_{k\to+\infty}\mathbb{E}_{0}[|X_{k}|^{4}]<+\infty$ (Theorem 4.6). Now, with $c_{n}$ being a constant depending on $n$ changing its values from line to line, due to
$\mathbb{E}_{k-1}\left[V_{k}^{2}\mathbbm{1}(|X_{k}|>\epsilon\sqrt{|M(t)|})\right]$ $=c_{n}\int_{x^{2}>\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}}x^{2}\left(1+\frac{x^{2}}{n-2}\right)^{-(n+1)/2}dx$
$\sim c_{n}\left(\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}\right)^{1-n/2},$
@@
$\qquad\sim c_{n}\epsilon^{2-n}\sum_{k\in M(t)}\left(\frac{\beta+\alpha X_{k-1}^{2}}{|M(t)|}\right)^{n/2},$
which tends to 0, $\mathbb{P}_{0}$-$a.s.$, since $\sum_{k\in M(t)}(\beta+\alpha X_{k-1}^{2})/|M(t)|$ is convergent to a strictly positive constant and $n\in\{5,6,\cdots\}$. This ensures the Lindeberg condition of Theorem 4.2. Next, to ensure the second condition of Theorem 4.6, observe that

### v2

@@ 5.2 Adaptive control variates
In this section, we examine the problem setting described in Examples 2.2 (c) and 3.2 (c). To recall it, suppose, for $\Psi:(0,1)^{d}\to\mathbb{R}$, we wish to evaluate the integral $\mu=\int_{(0,1)^{d}}\Psi({\bf u})d{\bf u}$, where the random element here is modeled via the uniform law on the unit hypercube $(0,1)^{d}$, which can be considered general enough thanks to the inverse transform random sampling method. Here, without affecting the value of the integral, one can introduce a parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ in a linear manner, as follows
@@
with ${\bm{\theta}}(0)=0_{d}$. By representing ${\bm{\theta}}(t)$ as a function of the batch index, we stress that it is updated only once for each batch. It clearly remains true that ${\bm{\theta}}(t)\to{\bm{\theta}}^{*}$, $\mathbb{P}_{0}$-$a.s.$ as $t\to+\infty$, just as in (5.6), because this is still nothing but standard iid Monte Carlo methods for the rightmost expression of (5.4). Here, the filtration $(\mathcal{F}_{k})_{k\in\mathbb{N}_{0}}$ can be set as $\mathcal{F}_{k}=\sigma(\mathcal{N})\vee\sigma(U_{1},\cdots,U_{k})$ for $k\in\mathbb{N}$, as well as the branched one $(\mathcal{F}_{k-1}^{\star})_{k\in M(\tau)}$ by $\mathcal{F}_{k-1}^{\star}=\mathcal{F}_{m(\tau_{\epsilon,\delta})}\vee\sigma(U_{m(\tau_{\epsilon,\delta}-1)+1}^{\star},\cdots,U_{k-1}^{\star})$ for all $k\in M(\tau_{\epsilon,\delta})$ on the iid regenerated batch $\{U_{k}^{\star}\}_{k\in M(\tau_{\epsilon,\delta})}$, where $\tau_{\epsilon,\delta}$ is a suitable stopping batch as in (5.1). Then, for every $t\in\mathbb{N}$ and $k\in M(t)$, the base formulation (Assumption 2.1) holds by setting $X_{k}=\Psi(U_{k})+\left\langle{\bm{\theta}}(t-1),U_{k}-\mathbbm{1}_{d}/2\right\rangle$, as
@@
It is worth noting that the proposed stopping rules for adaptive control variates have an interesting connection to averaged stochastic gradient descent (ASGD) [Bottou2018]. While conventional SGD implementations often rely on validation performance for early stopping [prechelt2012], monitoring averaged iterates, which helps reduce the effect of the noise term, has proven beneficial in practice. This suggests potential value in developing statistically driven stopping criteria that leverage averaging, as in our framework, particularly for validation-free scenarios [duvenaud2016]. ∎
@@ 6 Concluding remarks
We conclude this work by highlighting two promising directions for future research. The first involves improving the quality of the proposed stopping rules. Unlike stopping rules based on the classical central limit theorem (such as [BHST, FLYA2012]), a key challenge in enhancing the quality of stopping rules here arises from the lack of accurate finite-time approximations for the limiting law in the context of the martingale central limit theorem. Specifically, obtaining Berry-Esseen-type bounds based on higher-order moments, and moreover presenting them in an implementable form, remains out of reach. Although there is extensive literature on martingale central limit theorems and their rates of convergence - typically expressed as uniform bounds (such as [10.1214/21-AIHP1182, FAN20191028, 10.3150/07-BEJ6116]) and, in some cases, as non-uniform bounds (such as [EK1, pcc81, JOOS1991297]) - these results generally involve unspecifiable constants, limiting their practical applicability to the development of stopping rules, with an exception [sydney], which, to the best of our knowledge, provides the only implementable result of the uniform bound of Berry-Esseen type.
@@ 7 Proofs
Proof of Theorems 2.4 and 3.1.
Theorem 2.4 is elementary by the tower rule and the standing assumption $\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the unbiasedness, as well as due to, for instance, [W1971, VII.9] for the strong law of large numbers. For Theorem 3.1, the first result can be derived as
@@
where we have applied the $\mathcal{F}_{m(\tau)}$-measurability of the stopping batch $\tau$ for the first equality, the regeneration procedure (illustrated in Figure 1 and described alongside the statement of Theorem 3.1) for the second equality, and the tower rule with $\mathbb{E}^{\star}_{k-1}[X_{k}^{\star}]=\mathbb{E}_{k-1}[X_{k}]=\mu$ (Assumption 2.1) for the third and fourth equalities, due to $\mathcal{F}_{m(\tau-1)}\subseteq\mathcal{F}^{\star}_{k-1}$ for all $k\in M(\tau)=\{m(\tau-1)+1,\cdots,m(\tau)\}$. The second result is then straightforward due to $\mathcal{F}_{0}\subseteq\mathcal{F}_{m(\tau)}$. ∎
@@
A standard martingale central limit theorem (for instance, [ish, Theorem 2.5 (a)] with $\xi_{t,k}=(X_{k}-\mu)/\sqrt{|M(t)|v^{2}_{0}(t)}$) reads in the present framework as follows. If $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ and $(|M(t)|v^{2}_{0}(t))^{-1}\sum_{k\in M(t)}\mathbb{E}_{k-1}[|X_{k}-\mu|^{2}\mathbbm{1}(|X_{k}-\mu|>\epsilon\sqrt{|M(t)|}v_{0}(t))]\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}0$ for every $\epsilon>0$, then it holds that $\mathbb{P}_{0}(\sqrt{|M(t)|}(\mu(t)-\mu)/v_{0}(t)\leq x)\to\Phi(x)$ for all $x\in\mathbb{R}$, as $t\to+\infty$. This Lindeberg condition can be simplified as the one in Theorem 4.2, thanks to $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ and $\sup_{t\in\mathbb{N}}v^{2}_{0}(t)<+\infty$ (Assumption 4.1). The desired weak convergence (4.6) holds true by the Slutsky theorem due to the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$. ∎
@@
$\epsilon^{1/q}\tau_{\epsilon,\delta}^{0}$ $=\epsilon^{1/q}\min\left\{t\in\mathbb{N}:\,\mathbb{P}_{0}\left(\sqrt{|M(t)|}\frac{|\mu(t)-\mu|}{v_{0}(t)}>\sqrt{|M(t)|}\frac{\epsilon}{v_{0}(t)}\right)\leq\delta\right\}$
$\sim\inf\left\{t\in(0,+\infty):\,2\left(1-\Phi\left(\epsilon\sqrt{|M(\lfloor\epsilon^{-1/q}t\rfloor)|}/v(\lfloor\epsilon^{-1/q}t\rfloor)\right)\right)\leq\delta\right\},$
@@
For Proposition 4.5, the expression (4.7) can be rewritten as $\tau_{\epsilon,\delta}=\min\{t\in\mathbb{N}:\,\epsilon\sqrt{|M(t)|}/(v(t)+a(t))\geq\Phi^{-1}(1-\delta/2)\}$, which yields $\tau_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ as $\epsilon\to 0+$ or $\delta\to 0+$, due to $|M(t)|\to+\infty$ (Assumption 2.3), $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1). With the aid of this expression, we get $\epsilon\sqrt{|M(\tau_{\epsilon,\delta})|}/(v(\tau_{\epsilon,\delta})+a(\tau_{\epsilon,\delta}))\geq\Phi^{-1}(1-\delta/2)$ at the time $m(\tau_{\epsilon,\delta})$, and $\epsilon\sqrt{|M(\tau_{\epsilon,\delta}-1)|}/(v(\tau_{\epsilon,\delta}-1)+a(\tau_{\epsilon,\delta}-1))<\Phi^{-1}(1-\delta/2)$ at the time $m(\tau_{\epsilon,\delta}-1)$, which together give, in a similar manner to (7.2),
$\Phi^{-1}(1-\delta/2)\leq\frac{\epsilon\sqrt{|M(\tau_{\epsilon,\delta})|}}{v(\tau_{\epsilon,\delta})+a(\tau_{\epsilon,\delta})}$
$\,=\frac{\epsilon\sqrt{|M(\tau_{\epsilon,\delta}-1)|}}{v(\tau_{\epsilon,\delta}-1)+a(\tau_{\epsilon,\delta}-1)}\frac{\sqrt{|M(\tau_{\epsilon,\delta})|}}{\sqrt{|M(\tau_{\epsilon,\delta}-1)|}}\frac{v(\tau_{\epsilon,\delta}-1)+a(\tau_{\epsilon,\delta}-1)}{v(\tau_{\epsilon,\delta})+a(\tau_{\epsilon,\delta})}$
$\,<\Phi^{-1}(1-\delta/2)\frac{\sqrt{|M(\tau_{\epsilon,\delta})|}}{\sqrt{|M(\tau_{\epsilon,\delta}-1)|}}\frac{v(\tau_{\epsilon,\delta}-1)+a(\tau_{\epsilon,\delta}-1)}{v(\tau_{\epsilon,\delta})+a(\tau_{\epsilon,\delta})}\stackrel{{\scriptstyle a.s.}}{{\to}}\Phi^{-1}(1-\delta/2),$ (7.4)
as $\epsilon\to 0+$, due to the first result $\tau_{\epsilon,\delta}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$, as well as $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4), $v^{2}(t+1)/|M(t+1)|\sim v^{2}_{0}(t)/|M(t)|$ and $\inf_{t\in\mathbb{N}}v^{2}_{0}(t)>0$ (Assumption 4.1), and $\lim_{t\to+\infty}a(t)=0$. Hence, Part (ii) holds by the continuous mapping theorem since $\Phi$ is continuous on $\mathbb{R}$. With the additional condition, the first result of Part (iii) can be obtained by reformulating the equation (7.4). The second result of Part (iii) holds by the asymptotic equivalence (7.3) with the regenerated empirical mean $\mu_{\star}$ and the stopping rule $\tau_{\epsilon,\delta}^{a}$, which yields the desired weak convergence due to the martingale central limit theorem (4.6) with the aid of $v^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle a.s.}}{{\to}}1$ (Theorem 4.4) and $\tau_{\epsilon,\delta}^{a}\stackrel{{\scriptstyle a.s.}}{{\to}}+\infty$ (Proposition 4.5 (i)), where the two asymptotic equivalences above hold almost surely, respectively, due to the first result and the additional condition of Part (iii). ∎
@@ A.1 ARCH(1)
Recall that $X_{k}=(\beta+\alpha X_{k-1}^{2})^{1/2}V_{k}$ for all $k\in\mathbb{N}$ with $X_{0}=0$ and $\{V_{k}\}_{k\in\mathbb{N}}$ is a sequence of iid random variables with the $t$-distribution with $\nu$ degrees of freedom for some $\nu\in\{5,6,\cdots\}$, scaled in such a way to satisfy $\mathbb{E}_{0}[V_{1}^{2}]=1$, that is, the common probability density function is given by
$\frac{d}{dx}\mathbb{P}(V_{1}\leq x)=\frac{\Gamma((\nu+1)/2)}{\sqrt{(\nu-2)\pi}\Gamma(\nu/2)}\left(1+\frac{x^{2}}{\nu-2}\right)^{-(\nu+1)/2},\quad x\in\mathbb{R}.$ (A.1)
In particular, we have $\mathbb{E}_{0}[|V_{1}|^{4}]=3(\nu-2)/(\nu-4)$. It is trivial to derive that $\mathbb{E}_{k-1}[X_{k}]=(\beta+\alpha X_{k-1}^{2})^{1/2}\mathbb{E}_{k-1}[V_{k}]=0$ and $\mathbb{E}_{k-1}[X_{k}^{2}]=\beta+\alpha X_{k-1}^{2}$, due to $X_{k-1}\in\mathcal{F}_{k-1}$, $V_{k}\perp\mathcal{F}_{k-1}$, and $\mathbb{E}_{k-1}[V_{k}^{2}]={\rm Var}_{k-1}(V_{k})={\rm Var}_{0}(V_{k})=1$ for all $k\in\mathbb{N}$. Hence, we get $\mathbb{E}_{0}[X_{k}]=0$ (Assumption 2.1 with $\mu=0$) and ${\rm Var}_{0}(X_{k})=\mathbb{E}_{0}[X_{k}^{2}]=\beta+\alpha\mathbb{E}_{0}[X_{k-1}^{2}]=\cdots=\beta(1-\alpha^{k})/(1-\alpha)>0$, due to $\alpha\in(0,1)$, $\beta\in(0,+\infty)$ and $X_{0}=0$. Hence, the discrete-time stochastic process $\{X_{k}:\,k\in\mathbb{N}_{0}\}$ is asymptotically weakly stationary with mean $\mathbb{E}_{0}[X_{k}]=0$ and variance $\lim_{k\to+\infty}{\rm Var}_{0}(X_{k})=\beta/(1-\alpha)$, due to $\alpha\in(0,1)$. Hence, we get the convergence of the theoretical batch variance as
@@
One can show that all the relevant conditions hold true if the constant $\alpha$ and the degree of freedom $\nu$ satisfy $3\alpha^{2}(\nu-2)/(\nu-4)<1$. To verify the convergence of the conditional batch variance, observe first that for every $k\in\mathbb{N}$ and $m\in\{k+1,\cdots\}$,
@@
$=\frac{3(\nu-2)}{\nu-4}\left(\beta^{2}+2\beta\alpha\mathbb{E}_{0}\left[X_{k-1}^{2}\right]+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right)$
$\sim\frac{3(\nu-2)}{\nu-4}\left(\beta^{2}\frac{1+\alpha}{1-\alpha}+\alpha^{2}\mathbb{E}_{0}\left[|X_{k-1}|^{4}\right]\right).$
If $3\alpha^{2}(\nu-2)/(\nu-4)<1$, then we get
$\mathbb{E}_{0}\left[|X_{k}|^{4}\right]\to\frac{3\beta^{2}(1+\alpha)(\nu-2)}{(1-\alpha)(\nu-4-3\alpha^{2}(\nu-2))},$
${\rm Var}_{0}\left(X_{k}^{2}\right)\to\frac{3\beta^{2}(1+\alpha)(\nu-2)}{(1-\alpha)(\nu-4-3\alpha^{2}(\nu-2))}-\frac{\beta^{2}}{(1-\alpha)^{2}},$
@@
$\mathbb{P}_{0}$-$a.s.$, as $t\to+\infty$. With $v^{2}_{0}(t)\to\beta/(1-\alpha)$, the condition $v_{1}^{2}(t)/v^{2}_{0}(t)\stackrel{{\scriptstyle\mathbb{P}_{0}}}{{\to}}1$ (Theorem 4.4) holds true. In addition, the convergence of the sequence $\{\mathbb{E}_{0}[|X_{k}|^{4}]\}_{k\in\mathbb{N}}$ above implies the condition $\limsup_{k\to+\infty}\mathbb{E}_{0}[|X_{k}|^{4}]<+\infty$ (Theorem 4.6). Now, with $c_{\nu}$ being a constant depending on $\nu$ changing its values from line to line, due to
$\mathbb{E}_{k-1}\left[V_{k}^{2}\mathbbm{1}(|X_{k}|>\epsilon\sqrt{|M(t)|})\right]$ $=c_{\nu}\int_{x^{2}>\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}}x^{2}\left(1+\frac{x^{2}}{\nu-2}\right)^{-(\nu+1)/2}dx$
$\sim c_{\nu}\left(\frac{\epsilon^{2}|M(t)|}{\beta+\alpha X_{k-1}^{2}}\right)^{1-\nu/2},$
@@
$\qquad\sim c_{\nu}\epsilon^{2-n}\sum_{k\in M(t)}\left(\frac{\beta+\alpha X_{k-1}^{2}}{|M(t)|}\right)^{\nu/2},$
which tends to 0, $\mathbb{P}_{0}$-$a.s.$, since $\sum_{k\in M(t)}(\beta+\alpha X_{k-1}^{2})/|M(t)|$ is convergent to a strictly positive constant and $\nu\in\{5,6,\cdots\}$. This ensures the Lindeberg condition of Theorem 4.2. Next, to ensure the second condition of Theorem 4.6, observe that